OpenStudy (anonymous):

express in scientific notation. (6 X 10^3)(4 X 10^5)

OpenStudy (saifoo.khan):

\[(6 \times 10^3) (400 \times 10^3)\]

OpenStudy (saifoo.khan):

\[(6\times 400) \times 10^3\]

OpenStudy (saifoo.khan):

\[2400 \times 10^3\]

OpenStudy (saifoo.khan):

\[2.4 \times 10^6\]

OpenStudy (anonymous):

2400 * 10^3 = 2.4 * 10^6